# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Unknown Dynamics

### v1

@@ VI-B3 Analyses of the results
Figure 8: Simulation results of avoiding static and moving obstacles for four methods. The black arrow denotes the moving direction of the dynamic obstacle, and the black thin line is the desired path. Labels ‘EP’ and ‘SP’ denote the endpoint and start point, respectively.
TABLE III: Performance Evaluations of Fig. 8
Quantitative Metrics VF-LPC MPC-CBF LMPCC RHRL-KDP
$J_{\rm{MC}}$ $\mathbf{62.12}$ $69.34$ $80.79$ $121.20$
$\rm{CT}$ (s) $\mathbf{34.05}$ $40.00$ $37.55$    $42.78$
Length (m) $\mathbf{239.54}$ $278.27$ $262.83$ $283.99$
Aver. S.T. (s)    $\mathbf{0.05}$    $0.09$    $0.10$        $0.09$
@@
As seen in TABLE III, VF-LPC has the lowest motion control cost $J_{\rm{ML}}$. The computational cost of VF-LPC is lower than that of MPC-CBF, LMPCC and RHRL-KDP, and it generates the shortest path. As VF-LPC employs a linear time-invariant Koopman-based vehicle dynamic model to optimize the nonlinear optimal motion planning problem based on the scheme of Eb-RHRL, the computational time is the least. However, the MPC-based methods and RHRL-KDP require online solving of nonlinear optimization problems, resulting in a greater computational burden.
The adopted deep Koopman model is designed for real-time optimal planning and control of intelligent vehicles. Therefore, we demonstrate its effectiveness by showcasing the position and velocity error between the vehicle and the reference path during the task completion process. The simulation results are illustrated in Fig. 9. As depicted in the upper subfigure, within the obstacle-free region (00-$70$ m), the vehicle exhibits minimal velocity deviation, even when encountering obstacles with a reduction in velocity due to velocity planning. After avoiding the obstacle, the vehicle quickly regains its desired velocity. In the bottom subfigure, within the obstacle-free region, the vehicle demonstrates a minor tracking error (about $0.1\rm{m}$) in following the desired path. This observation underscores the VF-LPC’s commendable safety control performance.
Figure 9: The vehicle speed and distance error from the desired path of VF-LPC. The red dashed line at the upper panel represents the desired speed of $25$ km/h, and the blue solid line represents the speed of the vehicle. The bottom panel presents the distance error between the vehicle and the reference path.
@@ VI-C Model Uncertainties Learning
The uncertainty in the dynamics of a system negatively affects the planning and control performance of robots. To eliminate the influence, a quantified sparse GP [9] is utilized to compensate for the difference between the offline-trained deep Koopman model and the exact vehicle model. The simulated dynamical model can be derived from Eq. (43), revealing that the vehicle states $v_{x},v_{y},\omega$ and $\rho_{e}(x_{k})$ are primarily affected by dynamic parameters. Therefore, we further derive
@@
$\text{and}\ g(x,u)=g(v_{x},v_{y},\omega,\rho_{e}^{\top},a_{x},\delta_{f}):\mathbb{R}^{n_{\rho_{e}}+5}\rightarrow\mathbb{R}^{3}.$
@@
VI-D Real-World Experiments
@@
VI-D1 Multiple Static Obstacles Avoidance
@@
VI-D2 Moving Obstacles Avoidance
@@
From the overall tracking results in Fig. 12, the intelligent vehicle keeps safe distances from obstacles all the time and returns to the desired path smoothly. In addition, our method exhibits small lateral tracking errors and longitudinal velocity deviation Moreover, the maximum vehicle’s speed reaches $2.4\ \rm{m/s}$. Finally, it arrives at the ending point successfully and completes the safe tacking control task.
Figure 12: Avoiding multiple moving obstacles when tracking a desired path. The vehicles represent the intelligent vehicle in different time instants. Labels ‘EP’ and ‘SP’ denote the endpoint and start point, respectively.
@@ VII Conclusion and Future Work
This paper presents the Vector Field-guided Learning Predictive Control (VF-LPC) approach for mobile robots with safety guarantees, which offers a framework for integrated planning and tracking control of robots. VF-LPC incorporates robot dynamics constraints into the guiding vector field for safe robot maneuvering. This is a notable improvement over the existing composite vector fields. Also, the system dynamics are unknown and are learned using the deep Koopman operator. This learned model is then updated online by sparse GP to improve safety and control performance. In addition, VF-LPC guarantees safety by transforming the safety problem into the pursuit-evasion game between a pursuer and an evader; thus, a game-based exponential barrier function is proposed, and a nonlinear optimal control problem is optimized and solved within the receding-horizon RL framework. A rigorous theoretical analysis of Eb-RHRL is provided to guarantee online convergence of learning optimal policies. Extensive simulations and real-world experiments validate the superiority of the proposed VF-LPC approach.
Our work has several possible future directions. 1) VF-LPC is effective in collision-avoidance tests. Under an added prediction module, we believe that VF-LPC could be promising even in higher-speed tasks. 2) In unknown environments, it is still challenging to obtain prior information on obstacles, such as positions and contours. One may need to construct safe barrier functions in real-time using the sensor-perceived information. 3) The effectiveness of VF-LPC has been validated through real-world experiments. Future work could investigate and deploy VF-LPC on more complex dynamic systems.

### v2

@@ VI-B3 Analyses of the results
Figure 8: Simulation results of avoiding static and moving obstacles for four IMPC methods. The black arrow denotes the moving direction of the dynamic obstacle, and the black thin line is the desired path. Labels ‘EP’ and ‘SP’ denote the endpoint and start point, respectively.
TABLE III: Performance Evaluations of Integrated Motion Planning and Control Methods
Metrics VF-LPC MPC-CBF[42] LMPCC[43] RHRL-KDP[10] CFS[44]
$J_{\rm{MC}}$ $\mathbf{61.50}$ $76.76$ $79.09$ $155.98$ $105.03$
$L_{p}$ (m) $\mathbf{252.54}$ $287.45$ $266.76$ $307.29$ $273.71$
$t_{\mathrm{com}}$ (s) $\mathbf{38.40}$ $41.13$ $38.88$    $44.72$    $39.03$
$t_{\mathrm{cal}}$ (s) $\mathbf{0.006}$    $0.15$    $0.15$        $0.07$       $0.12$
@@
As seen in TABLE III, VF-LPC has the lowest motion control cost $J_{\rm{MC}}$. The computational cost of VF-LPC is lower than that of MPC-CBF, LMPCC, RHRL-KDP, and CFS, and it generates the shortest trajectory and has the least completion time. As VF-LPC employs a linear time-invariant Koopman-based vehicle dynamic model to optimize the nonlinear optimal motion planning problem based on the scheme of Eb-RHRL, the computational time is the least. However, the MPC-based methods and RHRL-KDP require online solving of nonlinear optimization problems, resulting in a computational burden.
@@ VI-C Model Uncertainties Learning
VI-D Validation on Unmanned Aerial Vehicles
In this subsection, we demonstrate the effectiveness of VF-LPC on motion planning tasks of a different class of mobile robots, i.e., quadrotor UAVs, which have intricate dynamics. We use a widely-used model [48] to simulate the dynamics.
Firstly, we conduct data-driven modeling of the quadrotor UAV using deep Koopman operators[21], followed by employing the VF-LPC to generate the robot’s maneuvers in a 3D environment, as illustrated in Fig. 10. The state variables of the training data consist of the XYZ coordinates and velocities ($v_{x},v_{y},v_{z}$) of the quadrotor UAV, with control variables of acceleration in the XYZ directions. A total of $148$ data, each comprising approximately $4500$ samples, were collected randomly in the state and control space for training and model validation purposes. The $6$-dimensional state vector $[X,Y,Z,v_{x},v_{y},v_{z}]^{\top}$ was processed through an encode layer, resulting in a final system model of $18$ state variables and $3$ control variables (i.e., $n_{\mathcal{K}}=18,n_{u}=3$). Then the obtained model is incorporated into VF-LPC, which focuses on planning 3-axis accelerations for attitude control. Not only the VF-LPC approach can avoid multiple static obstacles. Under suddenly appearing obstacles (see Fig. 10), the VF-LPC approach is also capable of handling them, and the average speed of the vehicle reaches $1\rm{m/s}$.
The results show that the VF-LPC approach can achieve near-optimal motion planning for robots with offline-trained data-driven dynamics.
Figure 10: Validating test on UAVs: Avoiding multiple static and suddenly appearing obstacles when tracking a desired trajectory.
VI-E Real-World Experiments
@@
VI-E1 Multiple Static Obstacles Avoidance
@@
VI-E2 Moving Obstacles Avoidance
@@
Figure 12: Avoiding multiple moving obstacles when tracking a desired path. Labels ‘EP’ and ‘SP’ denote the endpoint and start point, respectively.
From the overall tracking results in Fig. 12, the intelligent vehicle keeps safe distances from obstacles all the time and returns to the desired path smoothly. In addition, our method exhibits small lateral tracking errors and longitudinal velocity deviation Moreover, the maximum vehicle’s speed reaches $2.4\ \rm{m/s}$. Finally, it arrives at the ending point successfully and completes the motion planning and control task.
@@ VII Conclusion and Future Work
This paper presents the Vector Field-guided Learning Predictive Control (VF-LPC) approach for mobile robots with safety guarantees, which offers a framework for the Integrated Motion Planning and Control (IMPC) of robots. VF-LPC designs kinodynamic guiding vector field for safe robot maneuvering. This is a notable improvement over the existing composite vector fields. Also, the learned deep Koopman model is updated online by sparse GP to improve safety and control performance. It is then incorporated into LPC to solve nonlinear IMPC problems. Rigorous theoretical analysis is provided to witness the online learning convergence.
VF-LPC is evaluated against motion planning methods that employ MPC and RL in high-fidelity CarSim software. The results show that VF-LPC outperforms them under metrics of completion time, route length, and average solution time. To further show the effectiveness and generalization of our proposed approach, we carried out path-tracking control tests of a mobile vehicle on a racing road to validate the model uncertainties learning capability, and we also successfully implemented the approach on quadrotor UAVs. Finally, we conducted real-world experiments on a Hongqi E-HS3 vehicle.
Our work has several possible future directions. 1) With an additional prediction module, VF-LPC could be promising even in higher-speed tasks. 2) In unknown environments, one can construct safer barrier functions in real-time using the information perceived by onboard sensors.